OpenStudy (anonymous):

What is the slope-intercept form of 9x + 8y = 7?

OpenStudy (rogue):

Slope intercept form is \[y = mx + b\]What you are given is standard form \[ax + by = c\]\[9x + 8y = 7\]In order to get it to y-int form, first subtract 9x from both sides.\[8y = -9x + 7\]Then just divide both sides by 8.\[y = \frac {-9}{8}x + \frac {7}{8}\]Done.